OpenStudy (anonymous):

Algebra 2 help!!! 7.04 FLVS!!!!

OpenStudy (anonymous):

You are going to purchase a new car, but being a responsible consumer means doing a little bit of research first. First, you find the vehicle you are purchasing and its price. Vehicle: Chevy Volt Price: $39,145 Current interest rate: 3% 1. Using the function A(t)=P(1+rn)nt, create the function that represents your new car loan that is compounded monthly. The principle will be the price of the vehicle you selected, not how much you are putting down. 2. Being a smart financial planner, you want to figure out how many months it will be until your principal is paid down to $10,000.00. Solve for t and show all of your work. Note that t will be negative because the number of months will decrease the principal. 3.Lastly, you decide to keep track of your loan four times a month instead of monthly. Solve for the adjusted interest rate. Remember to use the formula A(t)=P[(1+rn)1c]cnt where c = 4. When solving for the adjusted interest rate, be sure to set it equal to 1+rn

OpenStudy (anonymous):

For #1 I got A(t) = 39,145(1+ 0.03/12)^(12t)

OpenStudy (mathmale):

Your response to the first question looks appriopriate. If I'm reading the second question properly, the main thing you'll need to do here is to solve the given equation A=P(1+m/12)^12t for t. Using logs, it's not all that hard to solve for t.

OpenStudy (mathmale):

Look at the equation A=P(1+m/12)^12t. It has the form a = b*c^d. Take either the common log or the natural log of both sides. Review the "rules of logs," if necessary, to ensure that you simplify as much as possible through the application of logs. These rules include log ab+ log a and log b log a/b = log a - log b log a^b = b*log a I need to get off the 'Net now, but hope this information will at least help you get started. I'll be back on Open Study later.

OpenStudy (anonymous):

ughhh ok haha thanks

OpenStudy (anonymous):

can you help @tHe_FiZiCx99

OpenStudy (the_fizicx99):

Just follow Mathmale's instructions, just use your given information ^_~

OpenStudy (anonymous):

I don't know how to do that honestly... Im in flvs and they give no examples could you please show it in steps if you don't mind :)

OpenStudy (anonymous):

what class???

OpenStudy (the_fizicx99):

Eh? Class? I'm taking multiple classes >.>

OpenStudy (anonymous):

oooo which ones?

OpenStudy (the_fizicx99):

o_o Ehhm, I'm technically in my first year of HS, but I'm finishing up with my Math classes. I'm trying to be in AP Calc before the year ends..

OpenStudy (anonymous):

im going into highschool and as a freshman im going to be in pre calc... so I have to take algebra 2 honors... anyway haha I really need to finish this assignment -.-" Could you help me?

OpenStudy (dumbcow):

i see a problem with this question, what is the monthly payment? the formula you are given is for compound interest growth. you want to know how many monthly payments it will take to get the balance down to 10,000

OpenStudy (anonymous):

im in an engineering program :/ @tHe_FiZiCx99 and ok I just don't understand it

OpenStudy (dumbcow):

hint \[a = b^t\] take log of both sides \[\ln a = \ln b^t\] \[\ln a = t \ln b\] \[\frac{\ln a}{\ln b} = t\]

OpenStudy (anonymous):

thanks @dumbcow but I don't even know where to start... I got number one but then I just have no idea what to do

OpenStudy (dumbcow):

10,000 = 39,145(1+ 0.03/12)^(12t) from my example a = 10,000/39,145 b = 1 + .03/12

OpenStudy (anonymous):

ok so you just put 10,000 in for A(t)

OpenStudy (anonymous):

then what? @dumbcow

OpenStudy (dumbcow):

\[12 t = \frac{\ln \frac{10,000}{39,145}}{\ln (1 + \frac{.03}{12})}\]

OpenStudy (anonymous):

ohhh ok and the decimal answer I think I need is 2222.48? @dumbcow

OpenStudy (precal):

ok did you remember to divide everything by 12, recall you are solving for t

OpenStudy (anonymous):

oh! I did not! umm so 185.21?

OpenStudy (anonymous):

@precal

OpenStudy (precal):

I got -53.1555

OpenStudy (anonymous):

how? @precal

OpenStudy (precal):

you need a calculator to get a decimal number and you need to divide 12 at the end since you have that 12t

OpenStudy (anonymous):

ok i don't understand this either. I have an identicle question and I got to number one and I have no idea what to do now.i need step by step instructions to be able to understand it. and flvs don't explain anything very well!!!!!:(